# Theory of two-component superfluidity of microcavity polaritons

### v1

@@ Appendix A Systematic interaction reduction
The Hamiltonian interaction terms as we left them were somewhat cumbersome. We outline in this section how we systematically reduce them to obtain the simplified form in Eq. (16). Let us begin with the 16 interaction terms
@@
To simplify and rearrange the 16 interaction terms, we will utilize commutation relations for the bosonic creation and annihilation operators Abrikosov , as well as symmetry properties of the terms. The commutation relations reflect the nature of bosons, which can occupy the same quantum state without any restriction. The basic commutation relations for bosonic operators are,
$[\hat{l}_{P},\hat{l}_{P^{\prime}}^{\dagger}]=\delta_{P,P^{\prime}}$ (50)
$[\hat{u}_{P},\hat{u}_{P^{\prime}}^{\dagger}]=\delta_{P,P^{\prime}}$ (51)
$[\hat{l}_{P},\hat{l}_{P^{\prime}}]=0,\quad[\hat{u}_{P},\hat{u}_{P^{\prime}}]=0$ (52)
@@
$[\hat{l}_{P},\hat{u}_{P^{\prime}}]=[\hat{l}_{P},\hat{u}_{P^{\prime}}^{\dagger}]=0$ (53)
@@
Appendix B Hopfield coefficients simplification
where $\varepsilon_{\text{LP}}(P)$ is the energy of the lower polariton branch, $\varepsilon_{\text{ph}}(P)$ is the photon dispersion, $\hbar\Omega_{R}$ is the Rabi splitting, representing the coupling strength between photons and excitons. Under certain conditions, it is possible to approximate $X_{P}\approx\frac{1}{\sqrt{2}}$, indicating an equal mixture of photonic and excitonic components in the lower polariton state. The approximation $X_{P}\approx\frac{1}{\sqrt{2}}$ holds when the term inside the square root in the definition of $X_{P}$ is equal to 1. Specifically, this occurs when,
$\varepsilon_{\text{LP}}(P)-\varepsilon_{\text{ph}}(P)=\hbar\Omega_{R}.$ (58)
This relationship implies that the energy difference between the lower polariton and the photon at momentum $P$ matches the Rabi coupling strength $\hbar\Omega_{R}$. To take advantage of this approximation the photon and exciton energies should be close to resonance at $P=0$, meaning that the detuning $\Delta_{0}$ is small and this is experimentally adjustable dufferwiel2015 . This allows the lower polariton energy $\varepsilon_{\text{LP}}(0)$ to be approximately halfway between the photon and exciton energies at $P=0$. The Rabi splitting $\Omega_{R}$ should be large enough relative to any small detuning between $\varepsilon_{\text{ph}}(P)$ and $\varepsilon_{\text{exc}}(P)$. This ensures significant mixing of photon and exciton states, leading to equal contributions in the lower polariton. Under these conditions, both $X_{P}$ and $C_{P}$ approach $\frac{1}{\sqrt{2}}$, indicating that the lower polariton state is an equal superposition of photon and exciton components. This simplification is useful because it allows us to treat the lower polariton as an idealized mixture, facilitating further calculations without loss of significant accuracy in this regime.
Using the approximation $X_{P}\approx\frac{1}{\sqrt{2}}$ is justified in many physical scenarios where the detuning is minimal, and the coupling is strong, as it reflects the experimentally relevant case of a balanced polariton. By assuming equal fractions, we simplify our analysis and capture the core physics of the coupled photon-exciton system in the microcavity. This approach is commonly used in polariton studies when the lower polariton branch is at resonance or near-resonance, making it both theoretically and experimentally meaningful to approximate $X_{P}\approx\frac{1}{\sqrt{2}}$.
Appendix C Expanded Polariton Spectrum and Effective Masses
@@
We plot the effective masses of the upper and lower polaritons as a function of the detuning and the Rabi splitting. In Fig. 20 below we plot these dependances.
@@
Appendix D Detailed Collective Excitation Spectrum for Two-Component Polaritons
The collective excitation spectrum is given by
$\Omega_{j,\mathbf{P}}=\sqrt{\frac{\omega_{1,\mathbf{P}}^{2}+\omega_{2,\mathbf{P}}^{2}+(-1)^{j-1}\sqrt{\beta}}{2}},$ (72)
where the single-branch frequencies are
$\omega_{j,\mathbf{P}}^{2}=\varepsilon_{j,\mathbf{P}}^{2}+2U_{q}n_{j,0}\varepsilon_{j,\mathbf{P}},$ (73)
with $j=1$ for lower polaritons (LP) and $j=2$ for upper polaritons (UP).
The energy spectrum $\varepsilon_{j,\mathbf{P}}$ is defined as,
$\varepsilon_{j,\mathbf{P}}=\frac{\varepsilon_{\text{ph},\mathbf{P}}+\varepsilon_{\text{exc},\mathbf{P}}}{2}\mp\frac{1}{2}\sqrt{\left(\varepsilon_{\text{ph},\mathbf{P}}-\varepsilon_{\text{exc},\mathbf{P}}\right)^{2}+4\hbar^{2}\Omega_{R}^{2}}.$ (74)
The photon dispersion is given by
$\varepsilon_{\text{ph},\mathbf{P}}=\frac{c}{n}\sqrt{\mathbf{P}^{2}+\left(\frac{\pi}{L_{c}}\right)^{2}}.$ (75)
For small $\mathbf{P}$, this simplifies to,
$\varepsilon_{\text{ph},\mathbf{P}}\approx\frac{c}{n}\left(\frac{\pi}{L_{c}}\right)+\frac{\hbar^{2}\mathbf{P}^{2}}{2m_{\text{ph}}},\quad m_{\text{ph}}=\frac{\hbar\pi}{cnL_{c}}.$ (76)
The exciton dispersion is given by
$\varepsilon_{\text{exc},\mathbf{P}}=E_{\text{band}}-E_{\text{binding}}+\frac{\hbar^{2}\mathbf{P}^{2}}{2m_{\text{exc}}}.$ (77)
The interaction strengths are,
$G_{jj}=U_{q}n_{j,0},$ (78)
where $n_{j,0}$ is the total polariton concentration for species $j$ and $U_{q}$ is the interaction constant. The cross-interaction term is given by:
$G_{12}=U_{q}\sqrt{n_{\text{LP},0}n_{\text{UP},0}}.$ (79)
The coupling and interaction terms contribute to $\beta$, defined as,
$\beta=\left(\varepsilon_{1,\mathbf{P}}^{2}+2U_{q}n_{\text{LP},0}\varepsilon_{1,\mathbf{P}}-\varepsilon_{2,\mathbf{P}}^{2}-2U_{q}n_{\text{UP},0}\varepsilon_{2,\mathbf{P}}\right)^{2}+16U_{q}^{2}n_{\text{LP},0}n_{\text{UP},0}\varepsilon_{1,\mathbf{P}}\varepsilon_{2,\mathbf{P}}.$ (80)
Substituting $\omega_{j,\mathbf{P}}^{2}=\varepsilon_{j,\mathbf{P}}^{2}+2U_{q}n_{j,0}\varepsilon_{j,\mathbf{P}}$, the collective excitation spectrum becomes,
$\Omega_{j,\mathbf{P}}=\sqrt{\frac{\varepsilon_{1,\mathbf{P}}^{2}+\varepsilon_{2,\mathbf{P}}^{2}+2U_{q}n_{\text{LP},0}\varepsilon_{1,\mathbf{P}}+2U_{q}n_{\text{UP},0}\varepsilon_{2,\mathbf{P}}+(-1)^{j-1}\sqrt{\beta}}{2}}.$ (81)

### v2

@@ Appendix A Systematic interaction reduction
We outline in this section how we systematically reduce Hamiltonian interaction terms to obtain the simplified form in Eq. (22). Let us begin with the 16 interaction terms
@@
To simplify and rearrange the 16 interaction terms, we will utilize commutation relations for the bosonic creation and annihilation operators AGD2012 , as well as symmetry properties of the terms. The commutation relations reflect the nature of bosons, which can occupy the same quantum state without any restriction. The basic commutation relations for bosonic operators are,
$\hat{l}_{P},\hat{l}_{P^{\prime}}^{\dagger}]=\delta_{P,P^{\prime}}$ (52)
$\hat{u}_{P},\hat{u}_{P^{\prime}}^{\dagger}]=\delta_{P,P^{\prime}}$ (53)
$\hat{l}_{P},\hat{l}_{P^{\prime}}]=0,\quad[\hat{u}_{P},\hat{u}_{P^{\prime}}]=0$ (54)
@@
$\hat{l}_{P},\hat{u}_{P^{\prime}}]=[\hat{l}_{P},\hat{u}_{P^{\prime}}^{\dagger}]=0$ (55)
@@
Appendix B Expanded Polariton Spectrum and Effective Masses
@@
We plot the effective masses of the upper and lower polaritons as a function of the detuning and the Rabi splitting. In Figure 8 below we plot these dependances.
@@
Appendix C Results for GaAs/AlGaAs quantum wells in a GaAs microcavity
We plot in this section all analogous results for GaAs/AlGaAs quantum wells in a GaAs microcavity. We begin by presenting plots of the sound velocity as a function of detuning and the Rabi splitting in Figure 9.
Figure 9: Sound velocity $c_{s}$ of a two-component polariton condensate: (a) dependence on detuning $\Delta_{0}$ at fixed $\Omega_{R}=15$ meV, and (b) dependence on Rabi splitting $\Omega_{R}$ at fixed $\Delta_{0}=20$ meV. The total polariton density is $n_{0}=1\times 10^{11}~\text{cm}^{-2}$.
Similarly we plot the sound velocity as a function of the total polariton density in Figure 10 below.
Figure 10: Sound velocity $c_{s}$ as a function of total polariton density $n_{0}$ for different population-split parameters $\alpha$. The Rabi splitting and detuning are fixed at $\Omega_{R}=15$ meV and $\Delta_{0}=10$ meV, respectively.
Next we present the results of the critical temperature $T_{c}$ as a function of the detuning and Rabi splitting, in Figure (11) below.
Figure 11: Critical temperature $T_{c}$ in a two-component polariton condensate. (a) Dependence on detuning $\Delta_{0}$ at fixed Rabi splitting $\Omega_{R}=15$ meV. (b) Dependence on Rabi splitting $\Omega_{R}$ at fixed detuning $\Delta_{0}=5$ meV. In both panels, the total polariton density is held constant at $n_{0}=1\times 10^{11}~\text{cm}^{-2}$, and results are shown for multiple values of the population–split parameter $\alpha$.
Afte that we present the results of the critical temperature $T_{c}$ as a function of the total polariton density in Figure (12) below.
Figure 12: Critical temperature $T_{c}$ in a two-component polariton condensate as a function of the total polariton density $n_{0}$ for multiple values of the condensate population–split parameter $\alpha$. The Rabi splitting is fixed at $\Omega_{R}=15$ meV and the detuning at $\Delta_{0}=5$ meV.
Finally we present the phenomenological inference plots for GaAs/AlGaAs below in Figure 13.
Figure 13: Condensate population–split parameter $\alpha$ in a two-component polariton condensate. (a) Dependence of $\alpha$ on the critical concentration $n_{c}$ at fixed system temperature $T=50$ K, Rabi splitting $\Omega_{R}=15$ meV, and detuning $\Delta_{0}=5$ meV. (b) Dependence of $\alpha$ on the critical temperature $T_{c}$ at fixed total polariton density $n_{0}=1\times 10^{11}~\text{cm}^{-2}$, Rabi splitting $\Omega_{R}=8$ meV, and detuning $\Delta_{0}=10$ meV.